how is 5 to 11 and 7 is to x how would you write this. example 5/11=7/x
Answer by Alan3354(69443)

how is 5 to 11 and 7 is to x how would you write this. example 5/11=7/x
It's 5:11::7:x
5 is to 11 as 7 is to x.
--------------
5/11=7/x gives the same result, tho.
5x = 77
x = 15.4